A tank of O2 used at home for a patient with pulmonary disease has a pressure of 220mmhg 2.0 moles of gas and a volume of 4.0l. What will the pressure be when the volume expands to 5.0L and 3.0 moles of O2 is added when the temperature does not change
1
Expert's answer
2020-06-22T09:16:23-0400

Initial

pressure of oxygen(P1)=220 mm of Hg

Pressure in atmosphere=220/760=22/76 atm

Volume of oxygen(V1)=4.0 L

Number of mole of oxygen(n1)=2.0

What will be the pressure(P2) when

Volume of oxygen(V2)=5.0 L

Number of mole(n2)=3.0

According to ideal gas equation

PV=nRT

Temperature is constant

PV/Rn=T

P1V1/Rn1=P2V2/Rn2

P2= P1V1Rn2/Rn1V2

= 22×4×3 ÷ 76×2×5

= 0.34736842105 atm

P2= 264 mm of Hg